# Spectral characterization of shadowing for linear operators on Hilbert spaces

### v1

@@ 1. Introduction and the main result
In finite-dimensional spaces and for normal operators on Hilbert spaces the above three notions are equivalent (see [4], [9], [15]). Hyperbolicity always implies uniform expansivity (see [9, Corollary 1]) and shadowing (see [15, Theorem 1]), but, in general, the converse implications are false. Examples of nonhyperbolic uniformly expansive operators were given by Eisenberg and Hedlund [9], while Bernardes et al. [4] have constructed nonhyperbolic operators with the shadowing property. For further examples of nonhyperbolic operators with the shadowing property, see  [2], [6] and Example 1.8.
@@
[9, Lemma 1] Let $X$ be a complex Banach space. An operator $T\in GL(X)$ is hyperbolic if and only if
@@
[10, Theorem 1] Let $X$ be a complex Banach space. An operator $T\in GL(X)$ is uniformly expansive if and only if
@@
It is easy to verify that $S=T^{*}$ so that $S^{*}=T$. Eisenberg and Hedlund [9, Example 4] have shown that
@@ 2. Preliminaries
[7, Chap. XI, Par. 1, Proposition 1.1, p. 347]  If $X$ is a complex Hilbert space, $T\in L(X)$ and $\lambda\in\mathbb{C}$, then the following statements are equivalent.
@@
The first part of the proposition, the equivalence of (i) and (ii), has been established by Cirilo et al. [6, Proposition 3]. The proof of the second part will be based on a surjectivity criterion for Banach space operators formulated in the next proposition. Recall that if $X$ is a complex Banach space and $T\in L(X)$, then the (normed-space) adjoint operator $T^{\prime}\in L(X^{\prime})$ is defined by $T^{\prime}(f)=f\circ T$ for $f\in X^{\prime}$, where $X^{\prime}$ denotes the dual space of $X$, the space of all bounded linear forms $f\colon X\rightarrow\mathbb{C}$.
@@
[14, Chap. II, Par. 9, Theorem 5, p. 86] If $X$ is complex Banach space and $T\in L(X)$, then the following statements are equivalent.
@@
As noted above, we need only to prove the equivalence of (ii) and (iii). It is known (see, e.g., Köthe [13, Chap. 5, Par. 26, Sec. 8, p. 359]) that the dual space of $l_{1}(X)$ is $l_{\infty}(X)$ if we identify each element $(y_{n})_{n\in\mathbb{Z}}\in l_{\infty}(X)$ with the bounded linear form $y\colon l_{1}(X)\rightarrow\mathbb{C}$ defined by
@@
Therefore, $x\in l_{\infty}(X)$ and $\mathcal{S}(x)=z$ with $\mathcal{S}$ as in (6). Since $z\in l_{\infty}(X)$ was arbitrary, this proves that $\mathcal{S}$ is surjective. According to [6, Proposition 3], the surjectivity of $\mathcal{S}$ implies that $T$ has the shadowing property. ∎
In the proof of Theorem 1.6, we shall use some facts about operator-valued holomorphic functions (see, e.g., [11, Chap. III, Sec. 3.11]). If $X$ is a complex Banach space and $R\colon A\rightarrow L(X)$ is an operator-valued holomorphic function defined on the annulus
@@
[12, Corollary 3.10] Suppose that $X$ is a complex Banach space and $G$ a domain (an open and connected set) in the complex plane. If $H\colon G\rightarrow L(X)$ is a holomorphic function such that $H(\lambda)$ is right invertible for every $\lambda\in G$, then there exists a holomorphic function $R\colon G\rightarrow L(X)$ satisfying $H(\lambda)R(\lambda)=I$ for all $\lambda\in G$.
@@ 3. Proof of the main result
Sufficiency.  Suppose that (3) is satisfied, i.e., $\lambda I-T$ is right invertible for every $\lambda\in\mathbb{C}$ with $|\lambda|=1$. Since the right resolvent set of $T$, the set of those $\lambda\in\mathbb{C}$ for which $\lambda I-T$ is right invertible, is an open subset of $\mathbb{C}$ (see, e.g., [7, Chap. VII, Par. 2, Theorem 2.2, p. 192]), this implies that $\lambda I-T$ is right invertible in a neighborhood of the unit circle $|\lambda|=1$. Thus, there exists $\epsilon>0$ such that the holomorphic operator-valued function $H(\lambda)=\lambda I-T$ is right invertible for every $\lambda$ from the open and connected set $A_{\epsilon}\subset\mathbb{C}$ given by
@@
Therefore, assumptions (7) and (8) of Proposition 2.5 are satisfied with the bounded linear operator $B=C_{-1}$. By the application of Proposition  2.5, we conclude that $T$ has the shadowing property.

### v2

@@ 1. Introduction and the main result
In finite-dimensional spaces and for normal operators on Hilbert spaces the above three notions are equivalent (see [5], [10], [16]). Hyperbolicity always implies uniform expansivity (see [10, Corollary 1]) and shadowing (see [16, Theorem 1]), but, in general, the converse implications are false. Examples of nonhyperbolic uniformly expansive operators were given by Eisenberg and Hedlund [10], while Bernardes et al. [5] have constructed nonhyperbolic operators with the shadowing property. For further examples of nonhyperbolic operators with the shadowing property, see  [2], [7] and Example 1.8.
@@
[10, Lemma 1] Let $X$ be a complex Banach space. An operator $T\in GL(X)$ is hyperbolic if and only if
@@
[11, Theorem 1] Let $X$ be a complex Banach space. An operator $T\in GL(X)$ is uniformly expansive if and only if
@@
It is easy to verify that $S=T^{*}$ so that $S^{*}=T$. Eisenberg and Hedlund [10, Example 4] have shown that
@@ 2. Preliminaries
[8, Chap. XI, Par. 1, Proposition 1.1, p. 347]  If $X$ is a complex Hilbert space, $T\in L(X)$ and $\lambda\in\mathbb{C}$, then the following statements are equivalent.
@@
The first part of the proposition, the equivalence of (i) and (ii), has been established by Cirilo et al. [7, Proposition 3]. The proof of the second part will be based on a surjectivity criterion for Banach space operators formulated in the next proposition. Recall that if $X$ is a complex Banach space and $T\in L(X)$, then the (normed-space) adjoint operator $T^{\prime}\in L(X^{\prime})$ is defined by $T^{\prime}(f)=f\circ T$ for $f\in X^{\prime}$, where $X^{\prime}$ denotes the dual space of $X$, the space of all bounded linear forms $f\colon X\rightarrow\mathbb{C}$.
@@
[15, Chap. II, Par. 9, Theorem 5, p. 86] If $X$ is complex Banach space and $T\in L(X)$, then the following statements are equivalent.
@@
As noted above, we need only to prove the equivalence of (ii) and (iii). It is known (see, e.g., Köthe [14, Chap. 5, Par. 26, Sec. 8, p. 359]) that the dual space of $l_{1}(X)$ is $l_{\infty}(X)$ if we identify each element $(y_{n})_{n\in\mathbb{Z}}\in l_{\infty}(X)$ with the bounded linear form $y\colon l_{1}(X)\rightarrow\mathbb{C}$ defined by
@@
The result of Proposition 2.5 is not really new. It can be deduced from a more general result by Backes, Dragičević and Singh [4, Theorem 3]. For completeness, we give an alternative short proof.
@@
Therefore, $x\in l_{\infty}(X)$ and $\mathcal{S}(x)=z$ with $\mathcal{S}$ as in (6). Since $z\in l_{\infty}(X)$ was arbitrary, this proves that $\mathcal{S}$ is surjective. According to [7, Proposition 3], the surjectivity of $\mathcal{S}$ implies that $T$ has the shadowing property. ∎
In the proof of Theorem 1.6, we shall use some facts about operator-valued holomorphic functions (see, e.g., [12, Chap. III, Sec. 3.11]). If $X$ is a complex Banach space and $R\colon A\rightarrow L(X)$ is an operator-valued holomorphic function defined on the annulus
@@
[13, Corollary 3.10] Suppose that $X$ is a complex Banach space and $G$ a domain (an open and connected set) in the complex plane. If $H\colon G\rightarrow L(X)$ is a holomorphic function such that $H(\lambda)$ is right invertible for every $\lambda\in G$, then there exists a holomorphic function $R\colon G\rightarrow L(X)$ satisfying $H(\lambda)R(\lambda)=I$ for all $\lambda\in G$.
@@ 3. Proof of the main result
Sufficiency.  Suppose that (3) is satisfied, i.e., $\lambda I-T$ is right invertible for every $\lambda\in\mathbb{C}$ with $|\lambda|=1$. Since the right resolvent set of $T$, the set of those $\lambda\in\mathbb{C}$ for which $\lambda I-T$ is right invertible, is an open subset of $\mathbb{C}$ (see, e.g., [8, Chap. VII, Par. 2, Theorem 2.2, p. 192]), this implies that $\lambda I-T$ is right invertible in a neighborhood of the unit circle $|\lambda|=1$. Thus, there exists $\epsilon>0$ such that the holomorphic operator-valued function $H(\lambda)=\lambda I-T$ is right invertible for every $\lambda$ from the open and connected set $A_{\epsilon}\subset\mathbb{C}$ given by
@@
Therefore, assumptions (7) and (8) of Proposition 2.5 are satisfied with the bounded linear operator $B=C_{-1}$. By the application of Proposition  2.5, we conclude that $T$ has the shadowing property. ∎